OpenStudy (aravindg):

USing f(x)=sinx/x prove that \[ \large \int\limits^{\pi/2}_{0}e^{-\sin x}dx<(\pi/2R)(1-e^{-R}) for R>0\]

OpenStudy (aravindg):

@ash2326 , @badreferences

OpenStudy (anonymous):

Alright, squeeze theorem.\[\int_0^{\frac\pi2}e^{-\sin x}\,dx\]Since the value for \(e^{-\sin x}\) fluctuates between \(e^x\) and \(\frac1{e^x}\), the area underneath the curve fluctuates as well. Can you guess where I want to go from here?

OpenStudy (anonymous):

"Fluctuates" is the wrong word. I meant to say "high" and "low" bounds. Sorry/

OpenStudy (anonymous):